1.     A company manufactures fuses. The percentage of non-defective fuses is 95.4%. A sample of 10 fuse was selected. Calculate the probability of selecting at least 3 defective fuses.     


1
Expert's answer
2022-03-01T09:48:48-0500

Let "X=" the number of defective fuses: "X\sim Bin (n, p)."

Given "n=10, p=1-0.954=0.046,q=0.954"


"P(X\ge 3)=1-P(X=0)-P(X=1)"

"-P(X=2)=1-\dbinom{10}{0}(0.046)^0(0.954)^{10-0}"

"-\dbinom{10}{1}(0.046)^1(0.954)^{10-1}"

"-\dbinom{10}{2}(0.046)^2(0.954)^{10-2}"

"\approx0.0091518085"